# A Watermark for Black-Box Language Models

### v1

@@ 4 Theory
Theorem 4.1 tells us that sampling tokens using our proposed scheme is, from a probabilistic perspective, indistinguishable from sampling from the underlying model, with the caveat that the unique seed values are conditionally independent given the counts of sequences. If we dismiss hash collisions as very low probability events, then since the key is fixed, this reduces to the assumption that unique $n$-grams across the sampled sequences are independent. How strong of an assumption this is depends on many factors such as $m$, the underlying distribution, and the counts $(c_{1},\dots,c_{j})$ themselves. One can construct cases where the assumption is reasonable and others where it is blatantly violated (e.g. if $n$-grams within a sequence are strongly correlated). One direction to making the assumption more palatable is to draw a fresh keys i.i.d. for each hash call. This would obviously destroy detectability. As a trade-off, one can leverage a set of secret keys (i.e. by drawing keys uniformly at random from a key set), which may reduce distortion, but will hurt detection as each key in the set needs to be tested against.
@@
No matter the choice of watermarking settings, assuming that the unique test $n$-grams are independent, we have the following equality on the false positive rate of Detect, using decision threshold $t$.
@@
This also holds for DetectRecursive if we further assume the $p$-values across secret keys are independent.
Selecting distinct independent secret keys $\{K_{1},\dots,K_{t}\}$ (and ignoring hash collisions that arise across calls to Detect within DetectRecursive), will help attain the necessary independence.
Although the alternative score distribution is generally intractable, with the strong assumption that there are no duplicate $n$-grams across the candidate sequences, then for a special choice of $F$, we can write the alternative in closed form and formulate the optimal detection test.
@@
For other choices of $F$, we can estimate $f_{1}$ via simulation. If we assume candidate sequences have the same length $k$ with no duplicate $n$-grams, then we can fill an $m\times k$ matrix with i.i.d. draws from $F$ and pick the first element of the row with the largest row-sum (among the $m$). We do this until we have sufficiently large (e.g. 10,000) samples from $f_{1}$. We apply a Gaussian kernel-density estimator where the bandwidth is chosen using Scott’s rule (Scott, 2015) to estimate $f_{1}(r)$ for test value $r$. Despite having $f_{0}$ in closed-form, for consistency, we can also estimate it non-parametrically by drawing from $F$.
@@ 5.1 Models, Datasets, and Hyperparameters
Models and Datasets. Our main model and dataset is Mistral-7B-instruct (Jiang et al., 2023) hosted on Huggingface [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 with bfloat16 quantization, and databricks-dolly-15k [^3] https://huggingface.co/datasets/databricks/databricks-dolly-15k (Conover et al., 2023), an open source dataset of instruction-following examples for brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization. We use prompts from the brainstorming, generation, open QA (i.e. general QA), and summarization categories, whose human responses are at least 50 tokens long (save one example, which was removed because the prompt was extremely long). For each of the 5233 total prompts, we generate two non-watermarked responses — a stochastic one using temperature 1, and the greedy / argmax decoding — along with a watermarked one for each scheme. We always force a minimum (maximum) of 250 (300) new tokens by disabling the stop token for the first 250 tokens, re-enabling it, and stopping the generation at 300, regardless of whether the stop token was encountered. To simulate real-world use, we de-tokenize the outputs to obtain plain text, and re-tokenize them during scoring. We study performance as a function of token length $T\leq 250$ by truncating to the first $T$ tokens.
For completeness, we also present the key results when Gemma-7B-instruct [^4] https://huggingface.co/google/gemma-7b-it with bfloat16 quantization is applied to the test split of eli5-category [^5] https://huggingface.co/datasets/rexarski/eli5_category. Prompts are formed by concatenating the the title and selftitle fields. Only examples with non-empty title and whose prompt contains a ? are kept — for a total of 4885 examples.
Hyperparameters. We consider the following choices of CDFs $F$ / $F_{k}$. (1) $F=U(0,1)$ and $F_{k}=\text{IrwinHall}(k)$. (2) $F=N(0,1)$ and $F_{k}=N(0,k)$. (3) $F=-\text{Gamma}(1/k,1)$ and $F_{k}=-\text{Exp}(1)$. (4) $F=\chi_{2}^{2}$ and $F_{k}=\chi_{2k}^{2}$.
@@ 5.2 Evaluation Metrics
We evaluate performance using three criteria.
@@
Distortion. Our scheme, along with most of the baselines, boasts a distortion-free property. This property comes with assumptions that are often violated in practice, for example by reuse of the secret key across watermarking calls. We quantify how faithful the watermarking procedure is to the underlying generative model by computing both the perplexity and likelihood of watermarked text under the generator (without watermarking). We include likelihood as the log-probabilities used in calculating perplexity can over-emphasize outliers.
@@ 5.3 Adversarial Attacks
Table 1: Main table of results, showing our black-box scheme (Flat) and its recursive variant (Rec.) for various $k$’s and $m$’s, along with baselines. PPL, WR and WR (3) refer to perplexity, win rate of a single response, and win rate of the best-of-3 responses respectively. pAUC is ROC-AUC up to max FPR of 1%. C and P stand for 10% corruption and paraphrasing attack. For paraphrasing, target lengths of {150, 200, 250} are used in calculating AUC and pAUC as performance is essentially random on shorter lengths. The standard errors are quite small and the maximum across rows is shown for each column. AUCs, pAUCS, and their standard errors are scaled by 100.
Figure 1: Performance of our flat scheme. Top: Detection AUC and pAUC with 1% max FPR for a range of target text lengths when there is no corruption. Bottom Left: AUC (mixed $T$’s) as a function of the average non-watermarked response entropy of the examples used in the calculation. $x$-coordinate $x$ corresponds to the bucket of examples whose entropy is between $[x-0.25,x]$ nats. Bottom Right: Effect of amount of random token corruption on AUC (mixed $T$’s).
@@ 5.4 Baselines
Aaronson (A). Aaronson (2023) computes a PRN for each token $i$ in the vocabulary as $u_{i}=U(0,1)[h(K|w|i)]$, where $w$ is the preceding $(n-1)$-gram, $K$ is the secret key and $h$ is a cryptographic hash. Token $i^{*}$ is selected, where $i^{*}=\operatorname{argmax}_{i}u_{i}^{1/p_{i}}$. At test time, $n$-grams $\{w_{i}\}_{i=1}^{T}$ are extracted from the query test and the detection score $s$ is $-\sum_{i=1}^{T}\log\left(1-R_{i}\right)$, where $R_{i}=U(0,1)\left[h(K|w_{i})\right]$. $n$ is set to 4. This choice strikes a good balance between generation quality / diversity and robustness to attacks. The scheme boasts a distortion-free property, but the generated text is a deterministic function of the prompt — i.e. only one generation is possible conditioned on a particular prompt.
@@
If $k=1$ and $F=U(0,1)$, then our watermark encoding can be viewed as a stochastic version of Aaronson (2023)’s. As $m\to\infty$, $c_{t}/m\overset{a.s.}{\to}p_{t}$, where $p_{t}$ and $c_{t}$ are the probability and observed occurrences of token $t$.
@@
Aaronson Corrected (AC). Aaronson (2023)’s detection score $s_{A}$ is not length-aware and consequently a single decision threshold across scores involving various lengths results in poor performance, as we later show. Observing that $s_{A}$ is a sum of log $p$-values, $s_{A}\sim\text{Gamma}(T,1)$, or equivalently, $2s_{A}\sim\chi_{2T}^{2}$ under the null that all test tokens are non-watermarked. We propose the new corrected detection score, $s=1-\text{Gamma}(T,1)(s_{A})=1-\chi_{2T}^{2}(2s_{A})$. For completeness we also experiment with a $p$-value computed in the way we do for our method — concretely as, $1-\text{IrwinHall}(T)\left(\sum_{i=1}^{T}R_{i}\right)$. Note that both transformations are monotonic so they have no effect on ROC-AUC when T is fixed.
@@
Kirchenbauer (KB). Kirchenbauer et al. (2023a) uses the $n$ previous tokens to pseudorandomly partition the vocabulary for the next token into two lists: a green list of size $\gamma V$ and a red list consisting of the remainder. A positive bias of $\delta$ is added to the logits of the green list tokens while those of the red list are left unchanged. This has the effect of modifying $p$ so that green list tokens are more probable. The score for a text consisting of $T$ tokens, $T_{g}$ of which were found to be green is, $s=(T_{g}-\gamma T)/\sqrt{T\gamma(1-\gamma)}$. We incorporate the latest updates to the algorithm, [^6] https://github.com/jwkirchenbauer/lm-watermarking such as including the current token in the $n$-gram and skipping duplicate $n$-grams at test time. We set $n=4$, $\gamma=0.25$, and $\delta\in\{0.5,1,2,3,4\}$.
@@
Kuditipudi (K). A drawback of using the last $n$ tokens as a basis for the PRF is that changing just one of them changes the output and hurts detection. Kuditipudi et al. (2023) addresses this limitation as follows. Consider a secret, finite ordered list of seeds of length $k$. Start watermarking by selecting a position in the seed list uniformly at random and apply the selection rule of Aaronson (2023) with the PRNG seeded to the current value. Advance to the next seed in the list (wrap-around if you are at the end) and repeat. Scoring is done by conducting a permutation test evaluating how compatible the query text is with the specific list of seeds used during encoding as opposed to any other random list of seeds of the same length. As the random starting position is not known during scoring, an alignment score based on the Levenshtein distance is given that considers alignments of various subsequences of the text and seeds. The proposed method is quite similar to Aaronson (2023) with the difference of using a fixed list of seeds (instead of input tokens to determine the seed) and using a permutation test for scoring. The upside is robustness to token substitution attacks; the downside is significantly higher computational cost for scoring. Larger $k$ offers more diversity and quality in generation but comes with costlier and weaker detection. The scheme is distortion-free. Following their work, we let $k=256$ and accelerate the permutation test by pre-computing 5000 reference values for the secret list using snippets from the train set of C4-realnewslike (Raffel et al., 2019) at the various target lengths we evaluate on.
@@ 5.5 Experimental Results
Table 1 shows results for baselines and our scheme using $F=U(0,1)$ and $p$-values for scoring, as detailed in Algorithms 1 and 2. For the recursive scheme, depth is $\text{lg}(m)$ (i.e. $m=2$ for each imaginary watermarker). Here, the negative class is non-watermarked argmax/greedy generations. Results for using stochastic (temperature 1) generations as the negative as well as the average likelihood scores are presented in Table 5 (Appendix); the trends remain the same. We summarize our observations on Mistral-7B-instruct on databricks-dolly-15k, which also hold for Gemma-7B-instruct on eli5-category (presented in the Appendix).
@@ 5.5.1 Overall performance of our flat and recursive schemes
Our scheme is a competitive option for white-box watermarking. Is it better to use our method or alternatives in the white-box setting? When $k=1,m=1024$, we are able to achieve better perplexity (2.61 vs. 2.81), better diversity (62.2% vs. 45.3% on best-of-3 win rates) and comparable detection performance than Aaronson (2023). Furthermore, it has better perplexity (2.61 vs. 3.55) and detection performance (97.7% vs. 87.8% AUC) than Kuditipudi et al. (2023). By cranking up $\delta$, Kirchenbauer et al. (2023a) can achieve strong detection but at the expense of perplexity. When matched on perplexity, we achieve better detection. For example, $\delta=0.5$ achieves 3.39 PPL and 73.2% AUC compared to our 2.61 PPL and 97.7% AUC. Gemma-7B-instruct on eli5-category with $k=1,m=1024$ outperforms Kuditipudi et al. (2023) and is on-par with Aaronson (2023) (see Appendix). Kirchenbauer et al. (2023a) with $\delta=0.5$ gives 1.649 PPL and 61.6% AUC whereas $k=1,m=1024$ gets us 1.610 PPL with 93.2% AUC and even 1.645 PPL with 89.7% AUC when $k=50,m=16$ (black-box).
@@ 5.5.3 Observations on detection
Length correction of Aaronson (2023) is crucial. Recall that the ROC-AUCs presented in Table 1 are computed over a pool of different lengths. Our $p$-value-based score for Aaronson (2023) improves detection significantly; for example, AUC goes from 71.7% to 97.9%. Table 7 (Appendix) shows that the sum-based $p$-value correction fares a bit worse, which was a little surprisingly given that this worked the best for our scheme, even for the $k=1$ case.
Sum-based $p$-values outperform Fisher ones. In Table 2 (Appendix), we observe that replacing our sum-based $p$-value (where $\mathcal{H}_{0}$ is that $\sum_{i}^{T}r_{i}\sim F_{T}$) by a Fisher combination of token-level $p$-values hurts detection performance. For example, when $k=1,m=2$, AUC degrades from 90.2% to 86.3%. Note that when $k=1$, this setting corresponds exactly to a stochastic version of Aaronson (2023).
@@
Detection performance improves sharply with test samples $T$. Figure 1 shows the effect of $T$ on AUC. We see sharp improvements w.r.t. to $T$, even when $k$ is large and $m$ is small, highlighting the power of more test samples to counteract a weaker watermark signal.
@@
Paraphrasing can be extremely effective at destroying watermarks. We observe that paraphrasing can effectively erase the watermark as detection performance for most methods is near random. Kuditipudi et al. (2023) and Kirchenbauer et al. (2023a) with large $\delta$ do better on AUC (but not so much on pAUC). Furthermore, in Figures 1 and 2 (Appendix) we observe that large amounts of random token corruption hurts our scheme and Aaronson (2023)’s more than it that of Kirchenbauer et al. (2023a) or Kuditipudi et al. (2023).

### v3

@@ 4 Theory
Theorem 4.1 tells us that sampling tokens using our proposed scheme is, from a probabilistic perspective, indistinguishable from sampling from the underlying model, with the caveat that the unique seed values are conditionally independent given the counts of sequences. If we dismiss hash collisions as very low probability events, then since the key is fixed, this reduces to the assumption that unique $n$-grams across the sampled sequences are independent. How strong of an assumption this is depends on many factors such as $m$, the underlying distribution, and the counts $(c_{1},\dots,c_{j})$ themselves. One can construct cases where the assumption is reasonable and others where it is blatantly violated (e.g. if $n$-grams within a sequence are strongly correlated). We study how well this assumption holds in practice in the Appendix. One direction to making the assumption more palatable is to draw a fresh keys i.i.d. for each hash call. This would obviously destroy detectability. As a trade-off, one can leverage a set of secret keys (i.e. by drawing keys uniformly at random from a key set), which may reduce distortion, but will hurt detection as each key in the set needs to be tested against.
@@
No matter the choice of watermarking settings, assuming that the unique test $n$-grams are independent, we have the following equality on the false positive rate of Detect (Algorithm 1), using decision threshold $t$.
@@
This also holds for DetectRecursive (Algorithm 2) if we further assume the $p$-values across secret keys are independent.
Selecting distinct independent secret keys $(K_{1},\dots,K_{t})$ (and ignoring hash collisions that arise across calls to Detect within DetectRecursive), will help attain the necessary independence.
Although the alternative score distribution is generally intractable, with the strong assumption that there are no duplicate $n$-grams across the candidate sequences, then for a special choice of $F$, we can write the alternative in closed form, formulate the optimal detection test and write down false-positive and false-negative rates exactly.
@@
For other choices of $F$, we can estimate $f_{1}$ via simulation. If we assume candidate sequences have the same length $k$ with no duplicate $n$-grams, then we can fill an $m\times k$ matrix with i.i.d. draws from $F$ and pick the first element of the row with the largest row-sum (among the $m$). We do this until we have sufficiently large (e.g. 10,000) samples from $f_{1}$. We apply a Gaussian kernel-density estimator where the bandwidth is chosen using Scott’s rule (scott2015multivariate) to estimate $f_{1}(r)$ for test value $r$. Despite having $f_{0}$ in closed-form, for consistency, we can also estimate it non-parametrically by drawing from $F$.
@@ 5.1 Models, Datasets, and Hyperparameters
Models and Datasets. Our main model and dataset is Mistral-7B-instruct (jiang2023mistral) hosted on Huggingface [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 with bfloat16 quantization, and databricks-dolly-15k [^3] https://huggingface.co/datasets/databricks/databricks-dolly-15k (DatabricksBlog2023DollyV2), an open source dataset of instruction-following examples for brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization. We use prompts from the brainstorming, generation, open QA (i.e. general QA), and summarization categories, whose human responses are at least 50 tokens long (save one example, which was removed because the prompt was extremely long). For each of the 5233 total prompts, we generate two non-watermarked responses — a stochastic one using temperature 1, and the greedy / argmax decoding — along with a watermarked one for each scheme. We always force a minimum (maximum) of 250 (300) new tokens by disabling the stop token for the first 250 tokens, re-enabling it, and stopping the generation at 300, regardless of whether the stop token was encountered. To simulate real-world use, we de-tokenize the outputs to obtain plain text, and re-tokenize them during scoring. We study performance as a function of token length $T\leq 250$ by truncating to the first $T$ tokens.
For completeness, we also present the key results when Gemma-7B-instruct (team2024gemma) [^4] https://huggingface.co/google/gemma-7b-it with bfloat16 quantization is applied to the test split of eli5-category [^5] https://huggingface.co/datasets/rexarski/eli5_category. Prompts are formed by concatenating the the title and selftitle fields. Only examples with non-empty title and whose prompt contains a ? are kept — for a total of 4885 examples.
Hyperparameters. We consider the following choices of CDFs $F$ / $F_{k}$. (1) $F=U(0,1)$ and $F_{k}=\text{IrwinHall}(k)$. (2) $F=N(0,1)$ and $F_{k}=N(0,k)$. (3) $F=-\text{Gamma}(1/k,1)$ and $F_{k}=-\text{Exp}(1)$. (4) $F=\chi_{2}^{2}$ and $F_{k}=\chi_{2k}^{2}$. We set $n=4$.
@@ 5.2 Evaluation Metrics
We evaluate performance using three criteria. For all metrics and methods, standard errors are computed by running watermark generation multiple times, keeping the secret keys fixed.
@@
Distortion. Our scheme, along with most of the baselines, boasts a distortion-free property. This property comes with assumptions that are often violated in practice, for example by reuse of the secret key across watermarking calls. We would like to quantify the amount of distortion actually induced by the watermark but doing so is no straightforward task.
A meaningful metric is the KL-divergence between the true non-watermarked next-token distribution and the watermarked one conditioned on the same context, where the former is used as the reference, for all contexts that occur in the non-watermarked generation. This is challenging because while the explicit watermarked distribution is easily derived for kirchenbauer2023watermark’s scheme which explicitly tweaks the probability of each token in the vocabulary, this is not the case for our scheme; we could estimate it via simulations at each step but this would be computationally expensive. As an alternative, we measure the perplexity of the watermarked text under the LLM sans watermarking. The caveat is that while perplexity is often used in the literature to measure text quality, the goal here is not necessarily to minimize it (which greedy decoding would do) but to verify that watermarking does not materially change perplexity levels compared to that of non-watermarked randomly sampled text. We also report likelihood as the log-probabilities used in calculating perplexity can over-emphasize outliers.
@@ 5.3 Adversarial Attacks
Table 1: Main table of results, showing our black-box scheme (Flat) and its recursive variant (Rec.) for various $k$’s and $m$’s, along with baselines. The $k=1$ setting can be executed very efficiently when white-box access is available. PPL, WR and WR (3) refer to perplexity, win rate of a single response, and win rate of the best-of-3 responses respectively. pAUC is ROC-AUC up to max FPR of 1%. C and P stand for 10% corruption and paraphrasing attack. For paraphrasing, target lengths of {150, 200, 250} are used in calculating AUC and pAUC as performance is essentially random on shorter lengths. The standard errors are quite small and the maximum across rows is shown for each column. AUCs, pAUCS, and their standard errors are scaled by 100.
Figure 1: Performance of our flat scheme. Top: Detection AUC and pAUC with 1% max FPR for a range of target text lengths when there is no corruption. Bottom Left: AUC (mixed $T$’s) as a function of the average non-watermarked response entropy of the examples used in the calculation. $x$-coordinate $x$ corresponds to the bucket of examples whose entropy is between [x0.25,x][x-0.25,x] nats. Bottom Right: Effect of amount of random token corruption on AUC (mixed $T$’s).
@@ 5.4 Baselines
Aaronson (A). aaronson computes a PRN for each token $i$ in the vocabulary as $u_{i}=U(0,1)[h(K|w|i)]$, where $w$ is the preceding (n1)(n-1)-gram, $K$ is the secret key and $h$ is a cryptographic hash. Token $i^{*}$ is selected, where $i^{*}=\operatorname{argmax}_{i}u_{i}^{1/p_{i}}$. At test time, $n$-grams $\{w_{i}\}_{i=1}^{T}$ are extracted from the query test and the detection score $s_{A}=-\sum_{i=1}^{T}\log\left(1-R_{i}\right)$, where $R_{i}=U(0,1)\left[h(K|w_{i})\right]$. $n$ is set to 4. This choice strikes a good balance between generation quality / diversity and robustness to attacks. The scheme boasts a distortion-free property, but the generated text is a deterministic function of the prompt — i.e. only one generation is possible conditioned on a particular prompt.
@@
If $k=1$ and $F=U(0,1)$, then our watermark encoding can be viewed as a stochastic version of aaronson’s. As $m\to\infty$, $c_{t}/m\overset{a.s.}{\to}p_{t}$, where $p_{t}$ and $c_{t}$ are the probability and observed occurrences of token $t$.
@@
Aaronson Corrected (AC). aaronson’s detection score $s_{A}$ is not length-aware and consequently a single decision threshold across scores involving various lengths results in poor performance, as we later show. Observing that $s_{A}$ is a sum of (negative) log $p$-values, $s_{A}\sim\text{Gamma}(T,1)$, or equivalently, $2s_{A}\sim\chi_{2T}^{2}$ under the null that all test tokens are non-watermarked. We propose the new corrected detection score, $s_{AC}=\text{Gamma}(T,1)(s_{A})=\chi_{2T}^{2}(2s_{A})$. For completeness we also experiment with a $p$-value computed in the way we do for our method — concretely as, $\text{IrwinHall}(T)\left(\sum_{i=1}^{T}R_{i}\right)$. Note that both transformations are monotonic so they have no effect on ROC-AUC when T is fixed.
@@
Kirchenbauer (KB). kirchenbauer2023watermark uses the $n$ previous tokens to pseudorandomly partition the vocabulary for the next token into two lists: a green list of size $\gamma V$ and a red list consisting of the remainder. A positive bias of $\delta$ is added to the logits of the green list tokens while those of the red list are left unchanged. This has the effect of modifying $p$ so that green list tokens are more probable. The score for a text consisting of $T$ tokens, $T_{g}$ of which were found to be green is, $s_{KB}=(T_{g}-\gamma T)/\sqrt{T\gamma(1-\gamma)}$. We incorporate the latest updates to the algorithm [^6] https://github.com/jwkirchenbauer/lm-watermarking such as including the current token in the $n$-gram and skipping duplicate $n$-grams at test time. We set $n=4$, $\gamma=0.25$, and $\delta\in\{0.5,1,2,3,4\}$.
@@
Kuditipudi (K). A drawback of using the last $n$ tokens as a basis for the PRF is that changing just one of them changes the output and hurts detection. kuditipudi2023robust addresses this limitation as follows. Consider a secret, finite ordered list of seeds of length $k$. Start watermarking by selecting a position in the seed list uniformly at random and apply the selection rule of aaronson with the PRNG seeded to the current value. Advance to the next seed in the list (wrap-around if you are at the end) and repeat. Scoring is done by conducting a permutation test evaluating how compatible the query text is with the specific list of seeds used during encoding as opposed to any other random list of seeds of the same length. As the random starting position is not known during scoring, an alignment score based on the Levenshtein distance is given that considers alignments of various subsequences of the text and seeds. The proposed method is quite similar to aaronson with the difference of using a fixed list of seeds (instead of input tokens to determine the seed) and using a permutation test for scoring. The upside is robustness to token substitution attacks; the downside is significantly higher computational cost for scoring. Larger $k$ offers more diversity and quality in generation but comes with costlier and weaker detection. The scheme is distortion-free. Following their work, we let $k=256$ and accelerate the permutation test by pre-computing 5000 reference values for the secret list using snippets from the train set of C4-realnewslike (2019t5) at the various target lengths we evaluate on.
@@ 5.5 Experimental Results
Table 1 shows results for baselines and our scheme using $F=U(0,1)$ and $p$-values for scoring, as detailed in Algorithms 1 and 2. For the recursive scheme, depth is $\text{lg}(m)$ (i.e. $m=2$ for each virtual watermarker). Here, the negative class is non-watermarked argmax/greedy generations. Results for using stochastic (temperature 1) generations as the negative as well as the average likelihood scores are presented in Table 5 (Appendix); the trends remain the same. We summarize our observations on Mistral-7B-instruct on databricks-dolly-15k, which also hold for Gemma-7B-instruct on eli5-category (presented in the Appendix).
@@ 5.5.1 Overall performance of our flat and recursive schemes
Our scheme is a competitive option for white-box watermarking. Is it better to use our method or alternatives in the white-box setting? When $k=1,m=1024$, we are able to achieve lower perplexity (2.61 vs. 2.81), better diversity (62.2% vs. 45.3% on best-of-3 win rates) than and comparable detection performance to aaronson. Furthermore, it has lower perplexity (2.61 vs. 3.55) and detection performance (97.7% vs. 87.8% AUC) than kuditipudi2023robust. By cranking up $\delta$, kirchenbauer2023watermark can achieve strong detection but at the expense of perplexity. When matched on perplexity, we achieve better detection. For example, $\delta=0.5$ achieves 3.39 PPL and 73.2% AUC compared to our 2.61 PPL and 97.7% AUC. Gemma-7B-instruct on eli5-category with $k=1,m=1024$ outperforms kuditipudi2023robust and is on-par with aaronson (see Appendix). kirchenbauer2023watermark with $\delta=0.5$ gives 1.649 PPL and 61.6% AUC whereas $k=1,m=1024$ gets us 1.610 PPL with 93.2% AUC and even 1.645 PPL with 89.7% AUC when $k=50,m=16$ (black-box).
@@ 5.5.3 Observations on detection
Length correction of aaronson is crucial. Recall that the ROC-AUCs presented in Table 1 are computed over a pool of different lengths. Our $p$-value-based score for aaronson improves detection significantly; for example, AUC goes from 71.7% to 97.9%. Table 7 (Appendix) shows that the sum-based $p$-value correction fares a bit worse, which was a little surprising given that this worked the best for our scheme, even for the $k=1$ case.
Sum-based $p$-values outperform Fisher ones. In Table 2 (Appendix), we observe that replacing our sum-based $p$-value (where $\mathcal{H}_{0}$ is that $\sum_{i}^{T}r_{i}\sim F_{T}$) by a Fisher combination of token-level $p$-values hurts detection performance. For example, when $k=1,m=2$, AUC degrades from 90.2% to 86.3%. Note that when $k=1$, this setting corresponds exactly to a stochastic version of aaronson.
@@
Detection performance improves sharply with test samples $T$. Figure 1 shows the effect of $T$ on AUC. We see sharp improvements with respect to $T$, even when $k$ is large and $m$ is small, highlighting the power of more test samples to counteract a weaker watermark signal.
@@
Paraphrasing can be extremely effective at destroying watermarks. We observe that paraphrasing can effectively erase the watermark as detection performance for most methods is near random. kuditipudi2023robust and kirchenbauer2023watermark with large $\delta$ do better on AUC (but not so much on pAUC). Furthermore, in Figures 1 and 2 (Appendix) we observe that large amounts of random token corruption hurts our scheme and aaronson’s more than it that of kirchenbauer2023watermark or kuditipudi2023robust.